OpenStudy (anonymous):

Wow. Here it goes. Any ideas about how to prove this? (1+cosx)/(1-cosx) + (1+sinx)/(1-sinx) = (2(cosx-cscx))/(cotx-cosx-cscx+1)

OpenStudy (anonymous):

\[\frac{ 1+\cos(x) }{ 1-\cos(x) }+\frac{ 1+\sin(x) }{ 1-\sin(x) }=2\frac{ \cos(x)-\csc(x) }{ \cot(x)-\cos(x)-\csc(x)+1 }\]

OpenStudy (anonymous):

just wanted to rewrite this first to see it better.

OpenStudy (anonymous):

Thanks! It does look better.

OpenStudy (anonymous):

Looks scary actually!

OpenStudy (anonymous):

Something tells me it looks more intimidating than it is.

OpenStudy (anonymous):

I hope you are right because my brain hurts at this point.

OpenStudy (anonymous):

I hope you're still there and the typing is just long???

OpenStudy (anonymous):

I think the problem may just want you to rearrange terms to prove they are equal but I'm not entirely sure.

OpenStudy (anonymous):

Should be a problem where we manipulate the left side so it becomes matched to the right.

OpenStudy (anonymous):

I didn't use any identities to show they are equal but I wonder if they are looking for a simpler answer using trig identities.

OpenStudy (anonymous):

ok well then all I did was algebra to find an answer.

OpenStudy (anonymous):

You make it sound easy. Will you share?

OpenStudy (anonymous):

\[\frac{1-\sin(x)+\cos(x)-\sin(x)\cos(x) }{ 1-\sin(x)-\cos(x)+\sin(x)\cos(x) }+\frac{ 1-\cos(x)+\sin(x)-\sin(x)\cos(x) }{ 1-\sin(x)-\cos(x)+\sin(x)\cos(x) }=2\frac{ \cos(x)-\frac{ 1 }{ \sin(x) } }{ \frac{ \cos(x)-1 }{ \sin(x) }-\cos(x)+1 }\]

OpenStudy (anonymous):

basically just tedious algebra, seems the post was cut off but its a lotta of distribution.

OpenStudy (anonymous):

this is the initial setup.

OpenStudy (anonymous):

Thanks. It is a bonus question anyway but I am desperate for every point right now. Math has always come easy for me till now. I study constantly and am still struggling.

OpenStudy (anonymous):

No problem, this seems to mostly be a problem to keep you busy rather than stretching your understanding.

OpenStudy (anonymous):

my final answer was \[\frac{ 2-2\sin(x)\cos(x) }{ (1-\cos(x))(1-\sin(x)) } = ""\]

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (anonymous):

No problem and good luck.

OpenStudy (anonymous):

I worked it through that far with your help, but can't figure out hos tomkeepmgoing to get it to equal the right side. Any more ideas?

OpenStudy (anonymous):

Wow. My typing shows I am getting really sleepy.